OpenStudy (isabel☺):

Downward forces of 200 N and 70 N, respectively, are required to keep a wooden block totally immersed in water and in oil. If the volume of the block is 0.1m^3 , det. the density of the oil and the density of the wood. Help me to find the solution! Pls. Ans. is 867.347 kg/m^3 , 795.5918 kg/m^3

OpenStudy (wolfe8):

Can you recall the formula for buoyant force being \[F=\rho Vg\]? What else do you know?

OpenStudy (isabel☺):

I don't exactly know those downward forces... can you give me FBD in this problem? :)

OpenStudy (wolfe8):

FBD? Well, based on one of Newton's laws, an action results in an equal reaction in magnitude. Buoyant force acts upwards, so it is equal to whatever force downwards required to fully immerse the block.

OpenStudy (isabel☺):

BF(buoyant force) = W in air - W in liquid

OpenStudy (isabel☺):

pls help me to get this problem uhmm

OpenStudy (isabel☺):

is it like

OpenStudy (isabel☺):

oops DRAW is not functioning

OpenStudy (isabel☺):

are the block, the oil and the water in one container? oil is at the top of the water? where is the block located?

OpenStudy (isabel☺):

@.Sam.

OpenStudy (anonymous):

Lost my connection and my answers. Know water density is 1000kg/m3. Here goes again: Block in water forces balance: 200N + (block density)(0.1)(9.8) - (1000)(0.1)(9.8)=0 block density calculates out to be 796 kg/m3. Block in oil forces balance:70N + (796)(0.1)(9.8) - (oil density)(0.98)=0 oil density =867 kg/m3.

OpenStudy (isabel☺):

what is 200N and 70N in there? is 200N = Weight in H2O? 70N = Weight in oil? if so what you did is : WinH2O+(block density)(Volume of block)(acce. due to gravity)=(H2O density)(Volume of block)(acce. due to gravity) why addition? isn't it BF(by H2O) = W (in air) - W(in H2O) = (block density)(volume of block)(acce. due to gravity) ? I agree with your answer..but PLS. help me to understand :D Thanks! @douglaswinslowcooper

OpenStudy (vincent-lyon.fr):

Free body diagrams are these:  W + T1 = UpThrust 1 W + T2 = UpThrust 2 Now, T1 and T2 are known; W depends on volume of block and density of wood; UpThrusts depend on volume of block and densities of water (case1) and oil (case2). Just convert these formulae using volume and densities, then plug-in the values. This should lead you to the answers.